### v1

# Simple solutions of the Yang–Baxter equation

Abstract.
We present a characterization of simple finite non-degenerate bijective set-theoretic solutions of the Yang–Baxter equation in terms of the algebraic structure of the associated permutation skew left braces. In particular, we prove that they need to have a unique minimal non-zero ideal and modulo this ideal one obtains a trivial skew left brace of cyclic type.
Key words and phrases:
Yang–Baxter equation, set-theoretic solution, simple solution, skew left brace.
2020 Mathematics Subject Classification:
Primary: 16T25; Secondary: 20N99
1. Introduction
The Yang–Baxter equation has its origins in the works of Yang [50] and Baxter [5]. Beyond its role in mathematical physics, this equation plays a crucial role in various branches of pure mathematics, particularly in quantum groups, Hopf algebras, and knot theory [7, 34, 39, 49, 32]. A comprehensive description of all solutions of the Yang–Baxter equation remains currently beyond reach. For this reason, Drinfeld in [25] suggested focusing on the set-theoretic version of this equation. A set-theoretic solution of the Yang–Baxter equation is a pair $(X,r)$, where $X$ is a non-empty set and $r\colon X\times X\to X\times X$ is a bijective map satisfying the braid equation
$(r\times{\operatorname{id}})({\operatorname{id}}\times r)(r\times{\operatorname{id}})=({\operatorname{id}}\times r)(r\times{\operatorname{id}})({\operatorname{id}}\times r)$
on $X\times X\times X$. The class of non-degenerate set-theoretic solutions holds particular importance. As usual we write
$r(x,y)=(\lambda_{x}(y),\rho_{y}(x))$
and thus non-degeneracy means that each $\lambda_{x}$ and $\rho_{y}$ is a bijection. Pioneer works [29, 26] brought algebraic tools to study a subclass of these solutions, the non-degenerate set-theoretic solutions that are involutive, i.e., also $r^{2}=\operatorname{id}$. Later in [38, 44], these methods were extended to the class of non-degenerate bijective solutions. Given a non-degenerate bijective set-theoretic solution to the Yang–Baxter $(X,r)$, the (left) derived solution of $(X,r)$, for which we will use the suggestive notation $(X,r^{\prime})$, is given by
$r^{\prime}(x,y)=(y,\sigma_{y}(x)),\quad\text{where}\quad\sigma_{y}(x)=\lambda_{y}\rho_{\lambda_{x}^{-1}(y)}(x).$
It is well-known that $(X,r^{\prime})$ also is a non-degenerate bijective set-theoretic solution, see for instance [36, Proposition 5.7]. It turns out that the derived solution plays a crucial role in the study of non-degenerate bijective solutions, for instance, being involutive or being injective for a non-degenerate bijective solution can be read from its derived solution [44, 37]. It is easy to verify that $X$ with respect to the binary operation $\triangleleft$ defined by $x\triangleleft y=\sigma_{y}(x)$ is a rack, i.e., $\sigma_{y}$ is bijective for all $y\in X$ and for any $x,y,z\in X$ it holds $(x\triangleleft y)\triangleleft z=(x\triangleleft z)\triangleleft(y\triangleleft z)$. A rack $(X,\triangleleft)$ is said to be a quandle if $x\triangleleft x=x$ for each $x\in X$. Any group $B$ with the operation given by $g\triangleleft h=h^{-1}gh$ defines a rack operation on $B$, which is clearly a quandle operation. Every subrack of such a quandle is called conjugation quandle. On the other hand a rack $(X,\triangleleft)$ defines a set-theoretic solution to the Yang–Baxter equation by $r(x,y)=(y,x\triangleleft y)$. This relation between solutions to the Yang–Baxter equation and racks has been noticed by Brieskorn in [8]. The notion of racks has been considered in the literature, mainly as a way to produce knot invariants [32, 24, 40, 8] and in the study of Nichols algebras over a group [1].
A characterization of arbitrary finite non-degenerate bijective solutions is presently beyond reach. Hence one needs to restrict the investigations to building blocks of solutions. Of course a first and obvious restriction is to deal with indecomposable solutions. Recall that a solution $(X,r)$ is said to be decomposable if $X=Y\cup Z$, a disjoint union of non-empty subsets, and $r$ restricted to $Y\times Y$ and $Z\times Z$ respectively defines a solution on $Y$ and $Z$. Second any non-degenerate bijective solution has an epimorphic image that is a simple solution, i.e. a solution $(X,r)$ of which any epimorphic image is either isomorphic to $(X,r)$ or a solution that $(Y,s)$ with $Y$ a singleton. Simple solutions are indecomposable and are the smallest non-trivial solutions and thus are the building blocks of all solutions. The definition of simple derived solutions appeared in [33], where a group theoretic characterization of simple quandles (i.e., derived solutions satisfying $r(x,x)=(x,x)$, i.e., being square-free) have been obtained. The definition of simple solution seems to have appeared first in [1], where the classification of simple racks has been improved given a concrete characterization of which groups give rise to a simple rack. Concerning involutive non-degenerate solutions, in [45], Vendramin introduced a notion of simple solutions (using the language of cycle sets). This definition of simplicity, however, coincides with the original one only within the class of indecomposable solutions. Cedó and Okniński [20, 16] (see also [19]) constructed finite involutive simple solutions. Recently in [21] they proved for a finite involutive simple solution $(X,r)$ that if $|X|$ is not a prime power then $|X|$ cannot be square-free. Moreover, if $|X|$ is a prime power $p^{n}$ then $n=1$, and there exist simple solutions of order $p$. As Joyce described simplicity for quandles in a group theoretic setting, Castelli in [11], characterized the simplicity of involutive non-degenerate solution in an algebraic setting, namely in the setting of skew left braces.
Recall, Rump [42] and Guarnieri and Vendramin[30] introduced skew left braces to provide an algebraic framework for studying arbitrary non-degenerate set-theoretic solutions of the Yang–Baxter equation. A skew left brace is a triple $(B,+,\circ)$ such that both $(B,+)$ and $(B,\circ)$ are groups and $a\circ(b+c)=a\circ b-a+a\circ c$ holds for all $a,b,c\in B$. One says that a skew left brace $B$ is of abelian type (it also is simply called a left brace) if $(B,+)$ is an abelian group. The terminology used in the theory of Hopf–Galois extensions suggests that the additive group determines the type of the skew left brace. For instance, if the additive group is cyclic, we say that the skew left brace is of cyclic type. Since their introduction, they have been widely studied [37, 19, 35, 43], and recently popularized in [46]. Every skew left brace $B$ naturally provides a set-theoretic solution $(B,r_{B})$ of the Yang–Baxter equation and skew left braces of abelian type provide involutive solutions (i.e solutions that are self-inverse). Specifically, $r_{B}\colon B\times B\to B\times B$ is given by $r_{B}(a,b)=(-a+a\circ b,(-a+a\circ b)^{-1}\circ a\circ b)$, where $c^{-1}$ denotes the inverse of $c\in B$ in the group $(B,\circ)$. Two fundamental skew left braces are associated with a finite non-degenerate bijective solution $(X,r)$: the structure skew left brace $G(X,r)$ and the permutation skew left brace $\mathcal{G}(X,r)$. In the next section we give rigorous definitions of these structures. Let us just mention that if the solution $(X,r)$ is finite so is the permutation skew left brace. Moreover, if two non-degenerate bijective solutions $(X,r)$ and $(Y,s)$ are isomorphic, then we have $G(X,r)\cong G(Y,s)$ as skew left braces, as well as the permutation groups $\mathcal{G}(X,r)\cong\mathcal{G}(Y,s)$. But the converse is not true. Given a finite skew left brace $B$, Bachiller, Cedó, Jespers, and Okniński [2, 3] provided a construction of all non-degenerate bijective set-theoretic solutions $(X,r)$ such that $\mathcal{G}(X,r)$ and $B$ are isomorphic as skew left braces. Providing more evidence that skew left braces are the right algebraic tool to investigate all non-degenerate bijective solutions of the Yang–Baxter equation. Within this framework, there is the need to investigate its building blocks, the simple skew left braces, i.e., skew left braces with no non-trivial homomorphic images. Note that in contrast with finite simple groups, it has been shown [16] that there is an abundance of finite simple skew left braces of abelian type [16, 17]. Moreover, simple skew left braces of abelian type have been constructed in [4]. Only recently, Byott [10] constructed a family of finite simple skew left braces that are not of abelian type. The first instances of this family can be found in the GAP package YangBaxter [47], i.e., SmallSkewbrace(12,22) and SmallSkewbrace(12,23).
Okniński and Cedó [20] proved that some finite simple skew left braces of abelian type determine an involutive simple solution $(X,r)$. They showed that finite simple braces which are additively generated by an orbit $X$ under the action of the permutation skew left brace $\mathcal{G}(X,r)$ yield simple non-degenerate involutive solutions on $X$. This connection between involutive simple solutions and simple skew left braces of abelian type recently has been refined by Castelli in [11] by showing that the simplicity of a finite non-degenerate involutive solution $(X,r)$ is described by the algebraic structure of its permutation skew left brace.
In this paper we determine when an arbitrary finite non-degenerate bijective solution is simple. As an application one recovers the results of Castelli, Cedó and Okniński and Joyce. Our main result can be summarized as follows, where we use the natural action of a skew brace $B$ on itself. Such an action can be restricted to ideals and subsets of $B$ that define subsolutions. The formulation also makes use of the ideal $B^{2}$. This is the ideal of $B$ generated by all the elements $-a+a\circ b-b$, for $a,b\in B$.
Theorem.
A finite non-degenerate bijective solution $(X,r)$ of the Yang–Baxter equation is simple if and only if it is one of the following types:
  1. (1)
    $(X,r)$ is a simple Lyubashenko solution.
  2. (2)
    $(X,r)$ is a simple derived solution.
  3. (3)
    $(X,r)$ is a subsolution of a solution associated with a skew left brace $B$ such that
    • $X$ generates $(B,+)$,
    • the ideal $B^{2}$ is the smallest non-zero ideal of $B$,
    • $B/B^{2}$ is a trivial skew left brace of cyclic type,
    • the action of the ideal $B^{2}$ on $X$ is transitive.
Recall that a solution $(X,r)$ is said to be Lyubashenko (see [25]) if $r(x,y)=(\lambda(y),\rho(x))$ for all $x,y\in X$, where $\lambda$ and $\rho$ are commuting permutations on $X$. Solutions of type (1) are completely characterized in Section 4, namely we prove that the cardinality of $X$ is a prime number and the group generated by $\lambda$ and $\rho$ is cyclic of order $p$. Solutions of type (2) are described in Section 3 and proofs of (2) and (3) also can be found in Section 3. As an application one obtains at once the classification of simple quandles by Joyce in [1] and the characterization of simple involutive solutions by Castelli in [11]. In (3) we obtain, in particular, that if the skew left brace $B$ is simple, then orbits (with respect to the natural action for $B$ on itself) generating $(B,+)$ provide simple solutions. In Section 5 we give several new classes of examples of simple solutions.
2. Preliminaries
Let $(X,r)$ be a finite non-degenerate bijective solution and $(X,r^{\prime})$ its (left) derived solution. We denote by
$M=M(X,r)=\langle X\mid x\circ y=\lambda_{x}(y)\circ\rho_{y}(x)\text{ for all }x,y\in X\rangle$
and
$A=A(X,r)=\langle X\mid x+y=y+\sigma_{y}(x)\text{ for all }x,y\in X\rangle$
the Yang–Baxter monoid (also known as structure monoid of $(X,r)$) and the left derived Yang–Baxter monoid of $(X,r)$ (also known as derived structure monoid of $(X,r)$), respectively (note that $A(X,r)=M(X,r^{\prime})$). Moreover,
$G=G(X,r)=\operatorname{gr}(X\mid x\circ y=\lambda_{x}(y)\circ\rho_{y}(x)\text{for all }x,y\in X)$
denotes the Yang–Baxter group (also known as structure group) of $(X,r)$ and
$A_{\operatorname{gr}}=A_{\operatorname{gr}}(X,r)=\operatorname{gr}(X\mid x+y=y+\sigma_{y}(x)\text{ for all }x,y\in X)$
denotes the left derived Yang–Baxter group (also known as left derived structure group) of $(X,r)$ (here $\langle X\mid S\rangle$ and $\operatorname{gr}(X\mid S)$ refer, respectively, to a monoid and group presentation with generators from a set $X$ and relations from a set $S$).
It was shown by Gateva-Ivanova and Majid [28, Theorem 3.6] that the maps $\lambda\colon X\to\operatorname{Sym}(X)\colon x\mapsto\lambda_{x}$ and $\rho\colon X\to\operatorname{Sym}(X)\colon x\mapsto\rho_{x}$ can be extended uniquely to $M=M(X,r)$ (for simplicity we use the same notation for the extension) so that one obtains a monoid morphism $\lambda\colon(M,\circ)\to\operatorname{Sym}(M)\colon a\mapsto\lambda_{a}$, and a monoid anti-morphism $\rho\colon(M,\circ)\to\operatorname{Sym}(M)\colon a\mapsto\rho_{a}$. Moreover, the map $r_{M}\colon M\times M\to M\times M$, defined as $r_{M}(a,b)=(\lambda_{a}(b),\rho_{b}(a))$, is a non-degenerate bijective solution of the Yang–Baxter equation. Furthermore, for $a,b,c\in M$, we have:
$a\circ b$ $=\lambda_{a}(b)\circ\rho_{b}(a),$ (1)
$\lambda_{a}(b\circ c)$ $=\lambda_{a}(b)\circ\lambda_{\rho_{b}(a)}(c),$ (2)
$\rho_{a}(b\circ c)$ $=\rho_{\lambda_{c}(a)}(b)\circ\rho_{a}(c).$ (3)
Also the map $\sigma\colon X\to\operatorname{Sym}(X)\colon x\mapsto\sigma_{x}$ can be extended to a monoid anti-morphism $\sigma\colon A\to\operatorname{Sym}(M)$ given by $\sigma_{b}(a)=\lambda_{b}\rho_{\lambda^{-1}_{a}(b)}(a)$. One obtains in the monoid $(A,+)$ that $a+b=b+\sigma_{b}(a)$ for all $a,b\in A$. It follows (see, e.g., [18, Proposition 3.2]) that there exists a bijective $1$-cocycle $\pi\colon M\to A$ (via the $\lambda$-map, i.e., $a+\lambda_{a}(b)=a\circ b$ for $a,b\in A$) such that $\pi(x)=x$ for all $x\in X$. So, we have a monoid embedding
$f\colon M\to A\rtimes{\operatorname{Im}(\lambda)}\colon m\mapsto(\pi(m),\lambda_{m}).$
Abusing notation, we will often identify $m$ with $f(m)$ or $\pi(m)$. Hence, in particular, we may also denote by $x_{i}$ the generators of $A$. So, we simply may write
$M=\{a=(a,\lambda_{a}):a\in A\}=\langle x=(x,\lambda_{x}):x\in X\rangle.$
Hence, in the terminology of [23], $(M,+,\circ,\lambda,\sigma)$ is a unital YB-semitruss, that is $(M,+)$ and $(M,\circ)$ are monoids (with the same identity $0=1$), $\lambda\colon(M,\circ)\to\operatorname{Aut}(M,+)$ is a monoid morphism with
$a\circ(b+c)=a\circ b+\lambda_{a}(c)\quad\text{and}\quad a+\lambda_{a}(b)=a\circ b,$
and thus
$\lambda_{a}\lambda_{b}=\lambda_{\lambda_{a}(b)}\lambda_{\rho_{b}(a)},$ (4)
$\sigma\colon(M,+)\to\operatorname{Aut}(M,+)$ is a monoid anti-morphism with
$a+b=b+\sigma_{b}(a)\quad\text{and}\quad\sigma_{\lambda_{a}(b)}\lambda_{a}(c)=\lambda_{a}\sigma_{b}(c),$ (5)
and thus also
$\sigma_{a}\sigma_{b}=\sigma_{\sigma_{a}(b)}\sigma_{a}$ (6)
for all $a,b,c\in M$. Note that for arbitrary left non-degenerate solutions one only requires that each $\sigma_{a}\in\operatorname{End}(M,+)$; but since we consider in this paper only non-degenerate bijective solutions all these maps are bijective. Because of the latter assumption, one obtains that the maps $\lambda$, $\rho$, and $\sigma$ extend to left and right actions on the structure group (see also [48]) and so that the equations (1), (2), and (3) remain valid in $G$; we use the same notation for these maps. Moreover, the bijective $1$-cocycle $M\to A$ naturally induces a bijective $1$-cocycle $G\to A_{\operatorname{gr}}$ (via the $\lambda$-map). Hence again we can identify $G$ with $A_{\operatorname{gr}}$ and thus $(G,+,\circ,\lambda,\sigma)$ is a YB-semitruss with $(G,+)$ and $(G,\circ)$ groups, in other words it is a skew left brace. Also,
$G=\{(a,\lambda_{a}):a\in A_{\operatorname{gr}}\}\subseteq A_{\operatorname{gr}}\rtimes{\operatorname{Im}(\lambda)}.$
So here we thus have
$\lambda_{a}(b)=-a+a\circ b\quad\text{and}\quad\sigma_{b}(a)=-b+a+b$
for all $a,b\in G$, and we have the associated solution $(G,r_{G})$ with $r_{G}(a,b)=(\lambda_{a}(b),\rho_{b}(a))$. Note that $X$ is naturally embedded in the structure monoid $M$. The natural map $i\colon X\to G$ is not necessarily injective (see, e.g., [37]) but it is a morphism of solutions. A solution $(X,r)$ is said to be injective if $i$ is injective. It is well known that involutive solutions are always injective.
In [15, Definition 1.5] the group
$\mathcal{G}_{\lambda,\rho}=\mathcal{G}_{\lambda,\rho}(X,r)=\operatorname{gr}((\lambda_{x},\rho_{x}^{-1}):x\in X)\subseteq\operatorname{Sym}(X)^{2},$
has been introduced and it is shown that $\mathcal{G}_{\lambda,\rho}$ is equipped with a skew left brace structure so that the natural map $G\to\mathcal{G}_{\lambda,\rho}$ is an epimorphism of skew left braces. We will consider an isomorphic copy of this skew left brace. For this we first need to prove some lemmas. The following lemma has been shown partially in [23, Proposition 4.3]. However an identity shown in [23, Lemma 3.3] gives an easier proof and a complete symmetric result. Put
$\mathfrak{q}\colon X\to X\colon x\mapsto\lambda_{x}^{-1}(x),$
known as the diagonal map. Recall that $\mathfrak{q}$ is a bijective map because $(X,r)$ is non-degenerate and bijective (see [23, Theorem 3.1] and [13]).
Lemma 2.1.
Let $a,b\in A$ and assume $\{\alpha,\beta,\gamma\}=\{\lambda,\rho,\sigma\}$. If $\alpha_{a}=\alpha_{b}$ and $\beta_{a}=\beta_{b}$ then also $\gamma_{a}=\gamma_{b}$. Furthermore, if $\alpha_{a}=\beta_{a}=\operatorname{id}$ then also $\gamma_{a}=\operatorname{id}$.
Proof.
From (5) and (4), and the definition of the $\sigma$-map, we get that for $a,b\in A$,
$\rho_{b}(a)=\lambda^{-1}_{\lambda_{a}(b)}\sigma_{\lambda_{a}(b)}(a)=\lambda^{-1}_{\lambda_{a}(b)}\lambda_{a}\sigma_{b}\lambda^{-1}_{a}(a)=\lambda_{\rho_{b}(a)}\lambda^{-1}_{b}\sigma_{b}\mathfrak{q}(a).$
Since, $\rho_{b}$ is bijective it yields $\lambda_{a}\lambda^{-1}_{b}\sigma_{b}\mathfrak{q}\rho^{-1}_{b}(a)=a$ for all $a\in A$. Hence,
$\mathfrak{q}(a)=\lambda^{-1}_{b}\sigma_{b}\mathfrak{q}\rho^{-1}_{b}(a),$
for all $a\in A$. So,
$\sigma_{b}=\lambda_{b}\mathfrak{q}\rho_{b}\mathfrak{q}^{-1}.$
The result now is obvious. ∎
As a consequence of the previous lemma we can now extend [23, Lemma 4.1] in the following sense. For completeness’ sake we include a proof.
Lemma 2.2.
Let $a,a^{\prime},b,b^{\prime}\in A$.
  1. (1)
    If $\sigma_{a}=\sigma_{a^{\prime}}$ and $\lambda_{a}=\lambda_{a^{\prime}}$ then $\lambda_{a\circ b}=\lambda_{a^{\prime}\circ b}$, $\rho_{b\circ a}=\rho_{b\circ a^{\prime}}$, $\sigma_{b\circ a}=\sigma_{b\circ a^{\prime}}$, and $\lambda_{a+b}=\lambda_{a^{\prime}+b}$, $\rho_{a+b}=\rho_{a^{\prime}+b}$, $\sigma_{a+b}=\sigma_{a^{\prime}+b}$.
  2. (2)
    If $\sigma_{b}=\sigma_{b^{\prime}}$ and $\lambda_{b}=\lambda_{b^{\prime}}$ then $\lambda_{a\circ b}=\lambda_{a\circ b^{\prime}}$, $\rho_{b\circ a}=\rho_{b^{\prime}\circ a}$, $\sigma_{b\circ a}=\sigma_{b^{\prime}\circ a}$, and $\lambda_{a+b}=\lambda_{a+b^{\prime}}$, $\rho_{a+b}=\rho_{a+b^{\prime}}$, $\sigma_{a+b}=\sigma_{a+b^{\prime}}$.
Proof.
(1) Assume $\sigma_{a}=\sigma_{a^{\prime}}$ and $\lambda_{a}=\lambda_{a^{\prime}}$. Because of Lemma 2.1 we have that also $\rho_{a}=\rho_{a^{\prime}}$. So we have that $\lambda_{a\circ b}=\lambda_{a}\lambda_{b}=\lambda_{a^{\prime}}\lambda_{b}=\lambda_{a^{\prime}\circ b}$ and similarly $\rho_{b\circ a}=\rho_{b\circ a^{\prime}}$. Further, $\sigma_{\lambda_{b}(a)}=\lambda_{b}\sigma_{a}\lambda^{-1}_{b}=\lambda_{b}\sigma_{a^{\prime}}\lambda^{-1}_{b}=\sigma_{\lambda_{b}(a^{\prime})}$ yields
$\sigma_{b\circ a}=\sigma_{b+\lambda_{b}(a)}=\sigma_{\lambda_{b}(a)}\sigma_{b}=\sigma_{\lambda_{b}(a^{\prime})}\sigma_{b}=\sigma_{b\circ a^{\prime}}.$
Now,
$\lambda_{a+b}=\lambda_{a\circ\lambda^{-1}_{a}(b)}=\lambda_{a}\lambda_{\lambda^{-1}_{a}(b)}=\lambda_{a^{\prime}}\lambda_{\lambda^{-1}_{a^{\prime}}(b)}=\lambda_{a^{\prime}+b},$
and similarly $\rho_{a+b}=\rho_{a^{\prime}+b}$. Moreover, $\sigma_{a+b}=\sigma_{b}\sigma_{a}=\sigma_{b}\sigma_{a^{\prime}}=\sigma_{a^{\prime}+b}$.
(2) Assume $\sigma_{b}=\sigma_{b^{\prime}}$ and $\lambda_{b}=\lambda_{b^{\prime}}$. Because of Lemma 2.1 we have that also $\rho_{b}=\rho_{b^{\prime}}$. So, we have $\lambda_{a\circ b}=\lambda_{a\circ b^{\prime}}$, $\rho_{b\circ a}=\rho_{b^{\prime}\circ a}$. Furthermore,
$\sigma_{b\circ a}=\sigma_{b+\lambda_{b}(a)}=\sigma_{\lambda_{b}(a)}\sigma_{b}=\sigma_{\lambda_{b^{\prime}}(a)}\sigma_{b^{\prime}}=\sigma_{b^{\prime}\circ a}.$
Moreover, $a+b=b+\sigma_{b}(a)=b\circ\lambda_{b}\sigma_{b}(a)$. Then
$\lambda_{a+b}=\lambda_{b\circ\lambda_{b}\sigma_{b}(a)}=\lambda_{b}\lambda_{\lambda_{b}\sigma_{b}(a)}=\lambda_{b^{\prime}}\lambda_{\lambda_{b^{\prime}}\sigma_{b^{\prime}}(a)}=\lambda_{a+b^{\prime}}.$
Similarly $\rho_{a+b}=\rho_{a+b^{\prime}}$. Finally, $\sigma_{a+b}=\sigma_{b}\sigma_{a}=\sigma_{b^{\prime}}\sigma_{a}=\sigma_{a+b^{\prime}}$. ∎
We now introduce the set
$\mathcal{G}_{\sigma,\lambda}=\mathcal{G}_{\sigma,\lambda}(X,r)=\{\operatorname{v}_{a}=(\sigma_{a}^{-1},\lambda_{a}):a\in A\}.$
Because of Lemma 2.2 we have the following well-defined binary operation on this set:
$\operatorname{v}_{a}+\operatorname{v}_{b}=\operatorname{v}_{a+b}\quad\text{and}\quad\operatorname{v}_{a}\circ\operatorname{v}_{b}=\operatorname{v}_{a\circ b}.$
Clearly $\mathcal{G}_{\sigma,\lambda}$ is additively (and multiplicatively) generated by the set $\{\operatorname{v}_{x}:x\in X\}$. As in [23, Section 4] one now obtains that $\mathcal{G}_{\sigma,\lambda}$ is a skew left brace for these operations with $\lambda$-map
$\lambda_{\operatorname{v}_{a}}(\operatorname{v}_{b})=\operatorname{v}_{\lambda_{a}(b)}$
and $\sigma$-map
$\sigma_{\operatorname{v}_{a}}(\operatorname{v}_{b})=\operatorname{v}_{\sigma_{a}(b)}.$
(Note that we use a $\lambda$ notation in several contexts, for example for defining set-theoretical solutions and on the skew left brace $\mathcal{G}_{\sigma,\lambda}$. It is clear from the context and the index used, which $\lambda$-map we consider. The same convention applies to the $\sigma$-map.)
Since the image of the following map
$\operatorname{v}\colon M=A\to\mathcal{G}_{\sigma,\lambda}\colon a\mapsto\operatorname{v}_{a}$
consists of invertible elements, both additively and multiplicatively, it induces an epimorphism of skew left braces (for which we use the same notation)
$\operatorname{v}\colon G\to\mathcal{G}_{\sigma,\lambda}\colon a\mapsto\operatorname{v}_{a}.$
The socle of the skew left brace $G$ is by definition
$\operatorname{Soc}(G)=\operatorname{Ker}(\operatorname{v})=\{a\in G:\lambda_{a}=\sigma_{a}=\operatorname{id}\}.$
It is an ideal of the skew left brace $G$. Recall that ideals of skew left braces are precisely the kernels of morphisms of skew left braces. It turns out that an additive subgroup $I$ of a skew left brace $(B,+,\circ)$ is an ideal if and only if it is $\lambda$-invariant, i.e., $\lambda_{a}(I)=I$ for all $a\in B$, $(I,+)$ is a normal subgroup of $(B,+)$ and $(I,\circ)$ is a normal subgroup of $(B,\circ)$. Equivalently, $I$ is $\lambda$-invariant, $(I,+)$ is a normal subgroup of $(B,+)$ and $i*a\in I$ for all $a\in B$ and $i\in I$. Recall that the operation *** is defined as follows:
$a*b=-a+a\circ b-b=\lambda_{a}(b)-b$
for all $a,b\in B$. An additive subgroup $I$ that is $\lambda$-invariant is said to be a left ideal of the skew left brace.
The solution $(X,r)$ is said to be irretractable [37] (see also, e.g., [26, 15]) if the restriction of the map $\operatorname{v}\colon M\to\mathcal{G}_{\sigma,\lambda}$ to $X$ is injective. In case $\operatorname{v}(X)$ is a singleton then the solution $(X,r)$ is a Lyubashenko solution.
Clearly, the restriction of the map $\operatorname{v}$ to $X$ is a morphism of solutions. Hence, if $(X,r)$ is simple then it follows that either $\operatorname{v}(X)$ is a singleton and we have a Lyubashenko solution, or the restriction of $\operatorname{v}$ to $X$ is an injective map, i.e., the solution $(X,r)$ is irretractable. Note that in the latter case $(X,r)$ is an injective solution and it is isomorphic with the solution induced on $\operatorname{v}(X)$. Furthermore, in the latter case $\mathcal{G}_{\sigma,\lambda}$ is additively (and also multiplicatively) generated by the set $\operatorname{v}(X)$.
Note that it also follows from Lemma 2.1 that the map $\mathcal{G}_{\sigma,\lambda}\to\mathcal{G}_{\lambda,\rho}$, defined as $\operatorname{v}_{a}\mapsto(\lambda_{a},\rho_{a}^{-1})$, is an isomorphism of skew left braces. From now on we work with the skew left brace $\mathcal{G}_{\sigma,\lambda}$ and we simply denote it as
$\mathcal{G}=\mathcal{G}_{\sigma,\lambda}.$
Let
$\mathcal{C}=\{\sigma_{a}:a\in A\}.$
It is a group for the multiplication
$\sigma_{a}\circ\sigma_{b}=\sigma_{a}\sigma_{b}=\sigma_{b+a}$
(note that the multiplication $\circ$ in $\mathcal{C}$ coincides, in fact, with the ordinary composition of maps) and we consider this as a trivial skew left brace, i.e., its additive structure is the same as its multiplicative structure. The multiplicative group $\mathcal{G}$ acts on $\mathcal{C}$ via the action $\cdot$ defined as follows:
${\operatorname{v}_{a}}\cdot{\sigma_{b}}=\sigma_{\lambda_{a}(b)}=\lambda_{a}\sigma_{b}\lambda_{a}^{-1}.$
Hence we may consider the semi-direct product of skew left braces
$\mathcal{S}=\mathcal{C}\rtimes\mathcal{G}.$
(For better readability we will denote elements of $\mathcal{S}$ as pairs $(\sigma_{a},\operatorname{v}_{b})$ with $a,b\in A$.) So, the additive operation is the componentwise addition and the multiplication is as follows:
$(\sigma_{a},\operatorname{v}_{b})\circ(\sigma_{c},\operatorname{v}_{d})=(\sigma_{a}\circ\sigma_{\lambda_{b}(c)},\operatorname{v}_{b\circ d})=(\sigma_{\lambda_{b}(c)+a},\operatorname{v}_{b\circ d}).$
Lemma 2.3.
The group $(\mathcal{S},\circ)$ acts on $X$ as follows:
$(\sigma_{a},\operatorname{v}_{b})\cdot x=\sigma_{a}\lambda_{b}(x).$
Proof.
If $a,b,c,d\in A$ and $x\in X$ then
$((\sigma_{a},\operatorname{v}_{b})\circ(\sigma_{c},\operatorname{v}_{d}))\cdot x$ $=(\sigma_{\lambda_{b}(c)+a},\operatorname{v}_{b\circ d})\cdot x$
$=\sigma_{\lambda_{b}(c)+a}\lambda_{b\circ d}(x)$
$=\sigma_{a}\sigma_{\lambda_{b}(c)}\lambda_{b}\lambda_{d}(x)$
and, by (5),
$(\sigma_{a},\operatorname{v}_{b})\cdot((\sigma_{c},\operatorname{v}_{d})\cdot x)$ $=(\sigma_{a},\operatorname{v}_{b})\cdot\sigma_{c}\lambda_{d}(x)$
$=\sigma_{a}\lambda_{b}\sigma_{c}\lambda_{d}(x)$
$=\sigma_{a}\sigma_{\lambda_{b}(c)}\lambda_{b}\lambda_{d}(x).$
As clearly $(\sigma_{0},\operatorname{v}_{0})\cdot x=x$, we see that indeed this is a left action. ∎
Note that the action of $(\sigma_{a},\operatorname{v}_{v})$ is denoted by $\theta_{(a,b)}$ in [14, 2].
3. Simple solutions versus skew left braces
In the skew left brace $\mathcal{S}=\mathcal{C}\rtimes\mathcal{G}$, as is common, we identify $\mathcal{C}$ and $\mathcal{G}$ as sub skew left braces. In particular, we simple write $(\sigma_{a},\operatorname{v}_{0})$ as $\sigma_{a}$ and $(\sigma_{0},\operatorname{v}_{a})$ as $\operatorname{v}_{a}$.
Lemma 3.1.
Assume that $(X,r)$ is a simple finite non-degenerate bijective solution of the Yang–Baxter equation with $|X|>2$. Let $\varphi\colon\mathcal{G}\to\mathcal{C}$ be a map defined as $\varphi(\operatorname{v}_{a})=\sigma_{a}^{-1}$ (note that $\varphi$ is a skew left brace epimorphism). If $I$ is a non-zero ideal of $\mathcal{G}$ such that $K=\varphi(I)\rtimes I$ is an ideal of the skew left brace $\mathcal{S}=\mathcal{C}\rtimes\mathcal{G}$ then the action of $(K,\circ)$ on $X$ is transitive.
Proof.
Let $a\in X$ and consider the $K$-orbit $K\cdot a$.
Claim 1. We have $\lambda_{x}(K\cdot a)=K\cdot\lambda_{x}(a)$ and $\sigma_{x}(K\cdot a)=K\cdot\sigma_{x}(a)$ for $x\in X$.
Indeed, let $(\sigma_{b},\operatorname{v}_{c})\in K$. Then
${\operatorname{v}_{x}}\circ(\sigma_{b},\operatorname{v}_{c})\circ\operatorname{v}_{x}^{-1}=(\sigma_{\lambda_{x}(b)},\operatorname{v}_{x\circ c})\circ\operatorname{v}_{x^{-1}}=(\sigma_{\lambda_{x}(b)},\operatorname{v}_{x\circ c\circ x^{-1}}),$
and thus by (5)
$\lambda_{x}((\sigma_{b},\operatorname{v}_{c})\cdot a)$ $=\lambda_{x}\sigma_{b}\lambda_{c}(a)$
$=\sigma_{\lambda_{x}(b)}\lambda_{x}\lambda_{c}(a)$
$=\sigma_{\lambda_{x}(b)}\lambda_{x\circ c\circ x^{-1}}\lambda_{x}(a)$
$=\left({\operatorname{v}_{x}}\circ(\sigma_{b},\operatorname{v}_{c})\circ\operatorname{v}_{x}^{-1}\right)\cdot\lambda_{x}(a)\in K\cdot\lambda_{x}(a),$
because, by assumption, $K$ is an ideal of $\mathcal{S}$. Similarly, since
$\sigma_{x}\circ(\sigma_{b},\operatorname{v}_{c})\circ\sigma_{x}^{-1}=(\sigma_{x}\sigma_{b},\operatorname{v}_{c})\circ\sigma_{-x}=(\sigma_{x}\sigma_{b}\sigma_{\lambda_{c}(-x)},\operatorname{v}_{c}),$
and because $K$ is an ideal of $\mathcal{S}$, we get, by (5),
$\sigma_{x}((\sigma_{b},\operatorname{v}_{c})\cdot a)$ $=\sigma_{x}\sigma_{b}\lambda_{c}(a)$
$=\sigma_{x}\sigma_{b}\lambda_{c}\sigma_{-x}\sigma_{x}(a)$
$=\sigma_{x}\sigma_{b}\sigma_{\lambda_{c}(-x)}\lambda_{c}\sigma_{x}(a)$
$=\left(\sigma_{x}\circ(\sigma_{b},\operatorname{v}_{c})\circ\sigma_{x}^{-1}\right)\cdot\sigma_{x}(a)\in K\cdot\sigma_{x}(a),$
as claimed.
Claim 2. Let $b\in A$ and $x\in X$.
  1. (1)
    If $\sigma_{b}\in K$ then $\lambda_{\sigma_{b}\cdot a}(x)=\lambda_{\sigma_{b}(a)}(x)\in K\cdot\lambda_{a}(x)$.
  2. (2)
    If $\operatorname{v}_{b}\in K$ then $\lambda_{\operatorname{v}_{b}\cdot a}(x)=\lambda_{\lambda_{b}(a)}(x)\in K\cdot\lambda_{a}(x)$.
  3. (3)
    If $\sigma_{b}\in K$ then $\sigma_{\sigma_{b}\cdot a}(x)=\sigma_{\sigma_{b}(a)}(x)\in K\cdot\sigma_{a}(x)$.
  4. (4)
    If $\operatorname{v}_{b}\in K$ then $\sigma_{\operatorname{v}_{b}\cdot a}(x)=\sigma_{\lambda_{b}(a)}(x)\in K\cdot\sigma_{a}(x)$.
Claim 2(1). Indeed, since $\sigma_{b}\in K$ we know that $\sigma_{b}=\varphi(\operatorname{v}_{c})$ for some $\operatorname{v}_{c}\in K$. So, $\sigma_{b}=\sigma_{c}^{-1}=\sigma_{-c}$. Since $\sigma_{-c}(a)=c+a-c$, we obtain $\sigma_{b}(a)=\sigma_{-c}(a)=c+a-c$, and thus
$\lambda_{\sigma_{b}(a)}(x)$ $={\operatorname{v}_{\sigma_{b}(a)}}\cdot x$
$={\operatorname{v}_{c+a-c}}\cdot x$
$=(\operatorname{v}_{c}+\operatorname{v}_{a}-\operatorname{v}_{c})\cdot x$
$=(\operatorname{v}_{a}-\operatorname{v}_{a}+\operatorname{v}_{c}+\operatorname{v}_{a}-\operatorname{v}_{c})\cdot x.$
Because $\operatorname{v}_{c}\in K$ and since $K$ is an ideal we know that $-\operatorname{v}_{a}+\operatorname{v}_{c}+\operatorname{v}_{a}\in K$, and thus also $k=-\operatorname{v}_{a}+\operatorname{v}_{c}+\operatorname{v}_{a}-\operatorname{v}_{c}\in K$. Consequently
$\lambda_{\sigma_{b}(a)}(x)=({\operatorname{v}_{a}}+k)\cdot x=\big{(}{\operatorname{v}_{a}}\circ\lambda^{-1}_{\operatorname{v}_{a}}(k)\big{)}\cdot x=\lambda_{a}\big{(}\lambda^{-1}_{\operatorname{v}_{a}}(k)\cdot x\big{)}.$
Again because $K$ is an ideal of $\mathcal{S}$ we have that $\lambda^{-1}_{\operatorname{v}_{a}}(k)\in K$. Hence $\lambda^{-1}_{\operatorname{v}_{a}}(k)\cdot x\in K\cdot x$ and thus, by Claim 1,
$\lambda_{\sigma_{b}(a)}(x)\in\lambda_{a}(K\cdot x)=K\cdot\lambda_{a}(x),$
as claimed.
Claim 2(2). We have
$\lambda_{\lambda_{b}(a)}(x)={\operatorname{v}_{\lambda_{b}(a)}}\cdot x=\lambda_{\operatorname{v}_{b}}(\operatorname{v}_{a})\cdot x=({\operatorname{v}_{b}}*{\operatorname{v}_{a}}+\operatorname{v}_{a})\cdot x.$
Because $K$ is an ideal of $\mathcal{S}$ and $\operatorname{v}_{b}\in K$ we have that $k=\operatorname{v}_{b}*\operatorname{v}_{a}\in K$. Hence
$\lambda_{\lambda_{b}(a)}(x)=\left({\operatorname{v}_{a}}+(-{\operatorname{v}_{a}}+k+{\operatorname{v}_{a}})\right)\cdot x$
Again, because $K$ is an ideal of $\mathcal{S}$, we have $l=-{\operatorname{v}_{a}}+k+{\operatorname{v}_{a}}\in K$. Hence $\lambda_{\operatorname{v}_{a}}^{-1}(l)\in K$, and then
$\lambda_{\lambda_{b}(a)}(x)$ $=({\operatorname{v}_{a}}+l)\cdot x$
$=\big{(}{\operatorname{v}_{a}}\circ\lambda_{\operatorname{v}_{a}}^{-1}(l)\big{)}\cdot x$
$=\lambda_{a}\big{(}\lambda_{\operatorname{v}_{a}}^{-1}(l)\cdot x\big{)}\in\lambda_{a}(K\cdot x)=K\cdot\lambda_{a}(x)$
by Claim 1.
Claim 2(3). We have
$\sigma_{\sigma_{b}(a)}(x)$ $=\sigma_{\sigma_{b}(a)}\cdot x$
$=\sigma_{-b+a+b}\cdot x$
$=(\sigma_{b}+\sigma_{a}-\sigma_{b})\cdot x$
$=(\sigma_{a}-\sigma_{a}+\sigma_{b}+\sigma_{a}-\sigma_{b})\cdot x.$
By assumption $\sigma_{a}\in K$ and thus also $\sigma_{b}+\sigma_{a}-\sigma_{b}\in K$. Therefore, $k=-\sigma_{a}+\sigma_{b}+\sigma_{a}-\sigma_{b}\in K$ and
$\sigma_{\sigma_{b}(a)}(x)=(\sigma_{a}+k)\cdot x=\sigma_{a}(k\cdot x).$
Hence, again by Claim 1, the claim follows.
Claim 2(4). We have
$\sigma_{\lambda_{b}(a)}(x)=\sigma_{\lambda_{b}(a)}\cdot x=\lambda_{\operatorname{v}_{b}}(\sigma_{a})\cdot x=({\operatorname{v}_{b}}*\sigma_{a}+\sigma_{a})\cdot x.$
Again because $K$ is an ideal of $\mathcal{S}$, we get that $k={\operatorname{v}_{b}}*\sigma_{a}\in K\cap\mathcal{C}$, and thus there exists $l\in K$ with $k+\sigma_{a}=\sigma_{a}+l$. Then
$\sigma_{\lambda_{b}(a)}(x)=(k+\sigma_{a})\cdot x=(\sigma_{a}+l)\cdot x=\sigma_{a}(l\cdot x).$
Hence, again by Claim 1, the claim follows.
Denote by $\mathcal{O}$ the set of $K$-orbits on $X$. From Claims 1 and 2 we obtain an epimorphism of solutions:
$X\to\mathcal{O}\colon x\mapsto K\cdot x.$
Since, by assumption, the solution $(X,r)$ is simple, we get that $|\mathcal{O}|=1$ or $|\mathcal{O}|=|X|$. The former means that the action of $K$ on $X$ is transitive, as desired. So, it remains to deal with the case $|\mathcal{O}|=|X|$. By assumption $I$ is a non-zero ideal of $\mathcal{G}$. So, it contains a non-zero element, say $\operatorname{v}_{a}$. Assume that $\lambda_{a}\neq\operatorname{id}$ and thus $\lambda_{a}(x)\neq x$ for some $x\in X$. Hence $K\cdot x$ contains $\lambda_{a}(x)$ and thus $K\cdot\lambda_{a}(x)=K\cdot x$, in contradiction with $|\mathcal{O}|=|X|$. Thus $K$ contains a non-zero element $\operatorname{v}_{a}$ with $\lambda_{a}=\operatorname{id}$. In particular, $\sigma_{a}\neq\operatorname{id}$. By the assumption on the ideal $K$ we know that $\sigma_{a}^{-1}\in K\cap\mathcal{C}$. Hence also $\sigma_{a}\in K\cap\mathcal{C}$. Again, let $x\in X$ be such that $\sigma_{a}(x)\neq x$. Then $\sigma_{a}(x)=\sigma_{a}\cdot x\in K\cdot x$ and thus $K\cdot\sigma_{a}(x)=K\cdot x$, again in contradiction with $|\mathcal{O}|=|X|$. This finishes the proof. ∎
Now we need one more step to prove our main result. This is based on an idea of Rump [41] and is used by Castelli [11]. For this we use the following. Let $V$ be an ideal of the skew left brace $\mathcal{S}$. Define the equivalence relation $\sim_{V}$ on $X$ as
$x\sim_{V}y\iff y=v\cdot x\text{ for some }v\in V,$
where $\cdot$ is the action of $\mathcal{S}$ on $X$ defined in Lemma 2.3.
Lemma 3.2.
Let $f\colon(X,r)\to(Y,s)$ be an epimorphism of finite non-degenerate bijective solutions of the Yang–Baxter equation. Define the equivalence relation $\sim_{f}$ on $X$ as follows:
$x\sim_{f}y\iff f(x)=f(y).$
Moreover, consider the additive subgroup of $\mathcal{G}$ defined as
$\mathcal{D}_{f}=\langle\operatorname{v}_{x}-\operatorname{v}_{y}:x,y\in X\text{ and }x\sim_{f}y\rangle_{+},$
called the displacement group of the morphism $f$ (the term used is suggested by the notion of displacement groups, sometimes also called transvection groups, within the theory of racks; see, e.g., [6, 32]). Let
$\mathcal{D}_{f}^{\prime}=\langle\sigma_{x}-\sigma_{y}:x,y\in X\text{ and }x\sim_{f}y\rangle_{+}=\langle\sigma_{x}^{-1}-\sigma_{y}^{-1}:x,y\in X\text{ and}x\sim_{f}y\rangle_{+}=\varphi(\mathcal{D}_{f})$
denote the displacement group of the morphism $(X,r^{\prime})\to(Y,s^{\prime})$ of derived solutions naturally induced by $f$ (where $\varphi$ is the map defined in Lemma 3.1). If $x\sim_{f}y$ for some $x,y\in X$ then:
  1. (1)
    $\alpha_{z}(x)\sim_{f}\alpha_{z}(y)$ for each $\alpha\in\{\lambda,\rho,\sigma\}$ and $z\in X$.
  2. (2)
    $\alpha_{x}(z)\sim_{f}\alpha_{y}(z)$ for each $\alpha\in\{\lambda,\rho,\sigma\}$ and $z\in X$.
  1. (3)
    $\mathcal{D}_{f}$ is an ideal of $\mathcal{G}$ and thus $\mathcal{C}\rtimes\mathcal{D}_{f}$ is an ideal of $\mathcal{S}=\mathcal{C}\rtimes\mathcal{G}$. Furthermore, $\mathcal{D}_{f}=\langle-\operatorname{v}_{x}+\operatorname{v}_{y}:x,y\in X\text{ and }x\sim_{f}y\rangle_{+}$.
  2. (4)
    $\mathbb{D}_{f}=\mathcal{D}_{f}^{\prime}\rtimes\mathcal{D}_{f}$ and $\mathcal{D}_{f}^{\prime}$ are ideals of $\mathcal{S}$.
  3. (5)
    $\sim_{\mathbb{D}_{f}}$ is a refinement of $\sim_{f}$, i.e., if $x\sim_{\mathbb{D}_{f}}y$ then $x\sim_{f}y$ for $x,y\in X$.
Proof.
The properties (1) and (2) follow straightforwardly from $f$ being an epimorphism of solutions.
(3) We know from (1) that $\mathcal{D}_{f}$ is a left ideal of $\mathcal{G}$. So, we need to prove that $(\mathcal{D}_{f},+)$ and $(\mathcal{D}_{f},\circ)$ are normal subgroups of the additive, respectively, multiplicative group $\mathcal{G}$. The latter is equivalent with proving that for $x\sim_{f}y$ we have $({\operatorname{v}_{x}}-{\operatorname{v}_{y}})*g\in\mathcal{D}_{f}$ for any $g\in\mathcal{G}$. It is sufficient to deal with $g=\operatorname{v}_{t}$ with $t\in X$. Indeed, once we have shown that $(\mathcal{D}_{f},+)$ is a normal subgroup, it is sufficient to note that in any skew left brace $B$ we have $a*(b+c)=a*b+b+a*c-b$ for all $a,b,c\in B$ (see, e.g., [31]).
We first prove that the subgroup $(\mathcal{D}_{f},+)$ is normal in $(\mathcal{G},+)$. For this let $x,y\in X$ with $x\sim_{f}y$. Then for any $t\in X$ we have $\sigma_{t}(x)\sim_{f}\sigma_{t}(y)$. Hence
$\sigma_{f(t)}(f(x))=f(\sigma_{t}(x))=f(\sigma_{t}(y))=\sigma_{f(t)}(f(y)).$
Now,
$-{\operatorname{v}_{t}}+({\operatorname{v}_{x}}-{\operatorname{v}_{y}})+\operatorname{v}_{t}=-{\operatorname{v}_{t}}+\operatorname{v}_{t}+\sigma_{\operatorname{v}_{t}}(\operatorname{v}_{x}-\operatorname{v}_{y})=\sigma_{\operatorname{v}_{t}}(\operatorname{v}_{x})-\sigma_{\operatorname{v}_{t}}(\operatorname{v}_{y})=\operatorname{v}_{\sigma_{t}(x)}-\operatorname{v}_{\sigma_{t}(y)}\in\mathcal{D}_{f}.$
Since this holds for all $t\in X$ it follows that indeed the additive subgroup is a normal subgroup.
Also note that $-\operatorname{v}_{y}+\operatorname{v}_{x}=-{\operatorname{v}_{x}}+(\operatorname{v}_{x}-\operatorname{v}_{y})+\operatorname{v}_{x}$ and thus $-\operatorname{v}_{y}+\operatorname{v}_{x}\in\mathcal{D}_{f}$. Hence $\langle-\operatorname{v}_{x}+\operatorname{v}_{y}:x,y\in X\text{ and }x\sim_{f}y\rangle_{+}\subseteq\mathcal{D}_{f}$. Similarly as above, one proves that $\langle-\operatorname{v}_{x}+\operatorname{v}_{y}:x,y\in X\text{ and }x\sim_{f}y\rangle_{+}$ is a normal subgroup, and that it contains $\mathcal{D}_{f}$. It follows that $\mathcal{D}_{f}=\langle-\operatorname{v}_{x}+\operatorname{v}_{y}:x,y\in X\text{ and }x\sim_{f}y\rangle_{+}$.
Because the diagonal map $\mathfrak{q}$ is bijective, we write $y=\lambda_{z}^{-1}(z)=\lambda_{z^{-1}}(z)$ for some $z\in X$. Hence $\operatorname{v}_{y}=\operatorname{v}_{\lambda_{{z^{-1}}}(z)}=\lambda_{\operatorname{v}_{z^{-1}}}(\operatorname{v}_{z})$. The latter means that $\operatorname{v}_{y}=-\operatorname{v}_{z^{-1}}+\operatorname{v}_{z^{-1}}\circ\operatorname{v}_{z}=-\operatorname{v}_{z^{-1}}$, as $\operatorname{v}_{z^{-1}}$ is the inverse of $\operatorname{v}_{z}$ in $\mathcal{G}$ for the multiplicative operation. Hence,
$-\operatorname{v}_{y}+\operatorname{v}_{x}=\operatorname{v}_{z^{-1}}+\operatorname{v}_{x}=\operatorname{v}_{z^{-1}}\circ\operatorname{v}_{\lambda_{z}(x)}.$ (7)
Consequently,
$(-\operatorname{v}_{y}+\operatorname{v}_{x})*\operatorname{v}_{t}=(\operatorname{v}_{z^{-1}}\circ\operatorname{v}_{\lambda_{z}(x)})*\operatorname{v}_{t}=\lambda_{\operatorname{v}_{z^{-1}}\circ\operatorname{v}_{\lambda_{z}(x)}}(\operatorname{v}_{t})-\operatorname{v}_{t}=\operatorname{v}_{\lambda_{z}^{-1}\lambda_{\lambda_{z}(x)}(t)}-\operatorname{v}_{t}.$
But we known from part (1) that $x\sim_{f}y$, i.e., $x\sim_{f}\lambda_{z}^{-1}(z)$, means that $\lambda_{z}(x)\sim_{f}z$ and thus this yields that $\lambda_{z}(t)\sim_{f}\lambda_{\lambda_{z}(x)}(t)$ for any $t\in X$ by part (2). Hence, again by part (1),
$t\sim_{f}\lambda_{z}^{-1}\lambda_{\lambda_{z}(x)}(t),$ (8)
and thus indeed $(-\operatorname{v}_{y}+\operatorname{v}_{x})*\operatorname{v}_{t}\in\mathcal{D}_{f}$.
(4) First we show that $\mathbb{D}_{f}$ is $\lambda$-invariant in $\mathcal{S}$. Since
$\lambda_{(\sigma_{a},\operatorname{v}_{b})}(\sigma_{c},\operatorname{v}_{d})$ $=-(\sigma_{a},\operatorname{v}_{b})+(\sigma_{a},\operatorname{v}_{b})\circ(\sigma_{c},\operatorname{v}_{d})$
$=(\sigma_{-a},\operatorname{v}_{-b})+(\sigma_{\lambda_{b}(c)+a},\operatorname{v}_{b\circ d})$
$=(\sigma_{\lambda_{b}(c)},\operatorname{v}_{\lambda_{b}(d)})$
and because $\lambda_{b}$ is an additive morphism, we get that
$\lambda_{(\sigma_{a},\operatorname{v}_{b})}(-\sigma_{x}+\sigma_{y},-\operatorname{v}_{k}+\operatorname{v}_{l})=(-\sigma_{\lambda_{b}(x)}+\sigma_{\lambda_{b}(y)},-\operatorname{v}_{\lambda_{b}(k)}+\operatorname{v}_{\lambda_{b}(l)}).$
Hence $\mathbb{D}_{f}$ and $\mathcal{D}_{f}^{\prime}$ are $\lambda$-invariant.
Next we show that $(\mathbb{D}_{f},+)$ and $(\mathcal{D}_{f}^{\prime},+)$ are normal subgroups of $(\mathcal{S},+)$. Clearly
$(\sigma_{c},\operatorname{v}_{d})+(\sigma_{a},\operatorname{v}_{b})-(\sigma_{c},\operatorname{v}_{d})=(\sigma_{-c+a+c},\operatorname{v}_{d+b-d})=(\sigma_{\sigma_{c}(a)},\operatorname{v}_{d+b-d}).$
So, with $a=y-x$, we get that the first coordinate is in $\mathcal{D}_{f}^{\prime}$ and if $b\in\mathcal{D}_{f}$, then by (3) the second coordinate belongs to $\mathcal{D}_{f}$. Moreover,
$(\sigma_{c},\operatorname{v}_{d})^{-1}\circ\sigma_{a}\circ(\sigma_{c},\operatorname{v}_{d})$ $=(\sigma_{\lambda^{-1}_{d}(-c)},\operatorname{v}_{d^{-1}})\circ(\sigma_{c+a},\operatorname{v}_{d})$
$=\sigma_{\lambda_{d}^{-1}(c+a)+\lambda_{d}^{-1}(-c)}$
$=\sigma_{\lambda_{d}^{-1}(-c)}\sigma_{\lambda_{d}^{-1}(a)}\sigma_{\lambda_{d}^{-1}(c)}.$
So, again, considering $a=y-x$, we obtain that $\mathcal{D}_{f}^{\prime}$ is also a multiplicative normal subgroup. In general, using (6), we have
$(\sigma_{c},\operatorname{v}_{d})^{-1}\circ(\sigma_{a},\operatorname{v}_{b})\circ(\sigma_{c},\operatorname{v}_{d})$ $=(\sigma_{\lambda^{-1}_{d}(-c)},\operatorname{v}_{d^{-1}})\circ(\sigma_{\lambda_{b}(c)+a},\operatorname{v}_{b\circ d})$
$=(\sigma_{\lambda_{d}^{-1}(\lambda_{b}(c)+a)+\lambda_{d}^{-1}(-c)},\operatorname{v}_{d^{-1}\circ b\circ d})$
$=(\sigma_{\lambda_{d}^{-1}(-c)}\sigma_{\lambda_{d}^{-1}(a)}\sigma_{\lambda_{d}^{-1}(\lambda_{b}(c))},\operatorname{v}_{d^{-1}\circ b\circ d})$
$=(\sigma_{u}\sigma_{\lambda_{d}^{-1}(a)}\sigma_{v},\operatorname{v}_{d^{-1}\circ b\circ d})$
$=(\sigma_{\sigma_{u}(\lambda_{d}^{-1}(a))}\sigma_{u}\sigma_{v},\operatorname{v}_{d^{-1}\circ b\circ d})$
$=(\sigma_{\sigma_{u}(\lambda_{d}^{-1}(a))}\sigma_{u}\sigma_{v},\operatorname{v}_{d}^{-1}\circ\operatorname{v}_{b}\circ\operatorname{v}_{d}),$
where $u=\lambda_{d}^{-1}(-c)$ and $v=\lambda_{d}^{-1}(\lambda_{b}(c))$. So, if $a=x-y$, then, as before, $\sigma_{u}\sigma_{v}\sigma_{\sigma_{v}(\lambda_{d}^{-1}(a))}\in\mathcal{D}_{f}^{\prime}$ and if $b$ is also a difference then we know from (3) that the second term is also in $\mathcal{D}_{f}$. This shows that $\mathbb{D}_{f}$ is a multiplicative normal subgroup of $\mathcal{S}$. Hence, $\mathbb{D}_{f}$ is indeed an ideal of the skew left brace $\mathcal{S}$.
(5) We need to prove that if $t=v\cdot z$, with $t,z\in X$ and $v\in\mathbb{D}_{f}$, then $f(t)=f(z)$. Now $v$ is a sum of elements of the form $\sigma_{-y+x}$ and $\operatorname{v}_{-x+y}$ with $x,y\in X$ such that $x\sim_{f}y$, i.e., $f(x)=f(y)$.
Assume first that $v=\operatorname{v}_{-y+x}$. Because of (7), $v=\operatorname{v}_{z}^{-1}\circ\operatorname{v}_{\lambda_{z}(x)}$. Hence $v\cdot t=\lambda_{z^{-1}\circ\lambda_{z}(x)}(t)$.
Let $x,y\in X$ be such that $\operatorname{v}_{x}\sim_{V_{f}}\operatorname{v}_{y}$. We want to prove that $x\sim_{f}y$, i.e., $f(x)=f(y)$. First, assume $v\in\mathbb{D}_{f}$ such that $x=v\cdot y$. Assume that $v=-\operatorname{v}_{k}+\operatorname{v}_{l}$ with $k,l\in X$ such that $f(k)=f(l)$ and, as the map $\mathfrak{q}$ is bijective, let $z\in X$ be such that $k=\lambda^{-1}_{z}(z)$. By (7) we know that $-\operatorname{v}_{k}+\operatorname{v}_{l}=\operatorname{v}_{z^{-1}}\circ\operatorname{v}_{\lambda_{z}(l)}$. And from the above we know that $v\cdot y=\lambda_{z^{-1}\circ\lambda_{z}(l)}(y)$. Hence,
$f(v\cdot y)=f(\lambda_{z^{-1}\circ\lambda_{z}(l)}(y))=\lambda_{f(z)^{-1}}\lambda_{\lambda_{f(z)}(f(l))}f(y).$
It follows from (8), where we use $t=y$, that $f(v\cdot y)=f(y)$. Second, assume $v=-\sigma_{k}+\sigma_{l}$ with $k,l\in X$ such that $f(k)=f(l)$. Then $v=\sigma_{l-k}$, and thus $x=v\cdot y=\sigma_{l-k}(y)$. Since $f\sigma_{z}=\sigma_{f(z)}f$ for each $z\in X$, we get
$f(x)=f(v\cdot y)=f(\sigma_{l-k}(y))=\sigma_{f(l)-f(k)}(f(y))=\sigma_{0}(f(y))=f(y).$
Consider now the general case, i.e., $v=v_{1}+\cdots+v_{n}$, with each term of the form $v_{i}=-\sigma_{x_{i}}+\sigma_{y_{i}}\in\mathcal{D}_{f}^{\prime}$ or $v_{i}=-\operatorname{v}_{x_{i}}+\operatorname{v}_{y_{i}}\in\mathcal{D}_{f}$. Then $v=w\circ w_{n}$ with $w=v_{1}+\cdots+v_{n-1}$ and $w_{n}$ an element of the form $-\sigma_{x}+\sigma_{y}$ or $-\operatorname{v}_{x}+\operatorname{v}_{y}\in\mathcal{D}_{f}$, with $x,y\in X$ satisfying $f(x)=f(y)$. Then $v\cdot t=w\cdot(w_{n}\cdot t)$. By the above and induction on $n$ it follows that that $v\cdot t\sim_{f}t$, as desired. ∎
To deal with the next step of the investigations we need to consider the structure group $G=G(X,r)$ of the simple finite non-degenerate bijective solution $(X,r)$. From the preliminaries, we know that the structure group $G$ is a skew left brace. Moreover, it may be identified with the regular subgroup $\{(a,\lambda_{a}):a\in A_{\operatorname{gr}}\}$ of the semi-direct product $A_{\operatorname{gr}}\rtimes\mathcal{G}_{\lambda}$. With this identification, the binary operations on $G$ are given by
$(a,\lambda_{a})+(b,\lambda_{b})$ $=(a+b,\lambda_{a+b}),$
$(a,\lambda_{a})\circ(b,\lambda_{b})$ $=(a+\lambda_{a}(b),\lambda_{a+\lambda_{a}(b)}).$
In particular, if $g=(a,\lambda_{a})\in G$ and $x\in X$ then
$\lambda_{g}(x,\lambda_{x})$ $=-(a,\lambda_{a})+(a,\lambda_{a})\circ(x,\lambda_{x})$
$=(-a,\lambda_{-a})+(a+\lambda_{a}(x),\lambda_{a+\lambda_{a}(x)})=(\lambda_{a}(x),\lambda_{\lambda_{a}(x)}).$
Now, let $I$ be an ideal of the skew left brace $G$ and assume there exists $g=(a,\lambda_{a})\in I$ with $\lambda_{a}\neq\operatorname{id}.$ By the previous
$\lambda_{g}(x,\lambda_{x})=(\lambda_{a}(x),\lambda_{\lambda_{a}(x)}).$
Now work in the skew left brace $G/I$ and consider the natural skew left brace epimorphism $\pi\colon G\to G/I$. Then
$\pi(\lambda_{a}(x),\lambda_{\lambda_{a}(x)})=\pi(\lambda_{g}(x,\lambda_{x}))=\lambda_{\pi(g)}\pi(x,\lambda_{x})=\lambda_{0}(\pi(x,\lambda_{x}))=\pi(x,\lambda_{x}).$
Thus
$(\lambda_{a}(x),\lambda_{\lambda_{a}(x)})-(x,\lambda_{x})\in I.$
As, by assumption, $\lambda_{a}\neq\operatorname{id}$ it follows that $(y,\lambda_{y})-(x,\lambda_{x})\in I$ for some distinct elements $x,y\in X$. Since, by assumption, $(X,r)$ is a simple solution, it follows that, $(y,\lambda_{y})-(x,\lambda_{x})\in I$ for all $x,y\in X$.
We get a similar property for the $\sigma$-map that determines the left derived group $A_{\operatorname{gr}}$. Indeed, let $I$ be an ideal of $G$ and suppose there exists $(a,\lambda_{a})$ with $\sigma_{a}\neq\operatorname{id}$. Let $x\in X$ be such that $\sigma_{a}(x)\neq x$ and consider again the canonical skew left brace epimorphism $\pi\colon G\to G/I$. Then
$\pi(\sigma_{a}(x),\lambda_{\sigma_{a}(x)})=\sigma_{\pi(a)}\pi(x,\lambda_{x})=\sigma_{0}\pi(x,\lambda_{x})=\pi(x,\lambda_{x}).$
Hence, $(\sigma_{a}(x),\lambda_{\sigma_{a}(x)})-(x,\lambda_{x})\in I$. As $(X,r)$ is simple, this implies $(y,\lambda_{y})-(x,\lambda_{x})\in I$ for all $x,y\in X$.
Lemma 3.3.
Assume $(X,r)$ is a simple finite non-degenerate bijective solution of the Yang–Baxter equation. Let $J$ be the ideal of $G$ generated by all the elements $(x,\lambda_{x})-(y,\lambda_{y})$ with $x,y\in X$. Let $I$ be an ideal of $G$. The following properties hold:
  1. (1)
    $J\subseteq I$ or $I\subseteq\operatorname{Soc}(G)$, the socle of $G$.
  2. (2)
    If $\operatorname{v}(J)\neq 0$ then $\operatorname{v}(J)$ is the smallest non-zero ideal of $\mathcal{G}$. In particular, if $\operatorname{v}(J)=\mathcal{G}$ then $\mathcal{G}$ is a simple skew left brace.
  3. (3)
    If $\operatorname{v}(J)=0$ then $(X,r)$ is a Lyubashenko solution.
Proof.
If $I$ contains an element $g=(a,\lambda_{a})$ with $\lambda_{a}\neq\operatorname{id}$ or $\rho_{a}\neq\operatorname{id}$ then we have already shown before the statement of the lemma that $J\subseteq I$. If $I$ does not contain such an element, then $I\subseteq\operatorname{Soc}(G)$ by Lemma 2.1 and thus (1) follows.
(2) Assume $\operatorname{v}(J)\neq 0$. Let $T$ be a non-zero ideal of $\mathcal{G}$. Then $\operatorname{v}^{-1}(T)$ is an ideal of $G$ properly containing the socle $\operatorname{Soc}(G)$. Thus, by (1), $J\subseteq\operatorname{v}^{-1}(T)$ and thus $\operatorname{v}(J)\subseteq T$. So, (2) follows.
(3) Assume $\operatorname{v}(J)=0$, i.e., for all $x,y\in X$ we have $\operatorname{v}_{x}=\operatorname{v}_{y}$. Hence $\sigma_{x}=\sigma_{y}$ and $\lambda_{x}=\lambda_{y}$ for all $x,y\in X$. So, $\rho_{x}=\rho_{y}$ for all $x,y\in X$ by Lemma 2.1, and thus $(X,r)$ is a Lyubashenko solution. ∎
We can now state the first part of our main theorem. Let $J$ be the ideal introduced in Lemma 3.3. By
$\mathcal{D}=\mathcal{D}(X,r)=\operatorname{v}(J)\quad\text{and}\quad\mathbb{D}=\mathbb{D}(X,r)=\varphi({\operatorname{v}(J)})\rtimes\operatorname{v}(J)$
we mean $\mathcal{D}_{f}$, respectively $\mathbb{D}_{f}$, for a morphism $f\colon(X,r)\to(Y,s)$ of solutions with $Y$ a singleton.
Theorem 3.4.
Let $(X,r)$ be a finite non-degenerate bijective solution of the Yang–Baxter equation with $|X|>1$. If $(X,r)$ is not a Lyubashenko solution then $(X,r)$ is simple if and only if $(X,r)$ is irretractable, the ideal $\mathcal{D}$ is the smallest non-zero ideal of $\mathcal{G}$ and the group $\mathbb{D}$ acts transitively on $X$.
Proof.
In the previous lemmas we have proved the necessity of the conditions. So, it suffices to prove the sufficiency. We prove this by contradiction. So assume $(X,r)$ is not simple, i.e., there exists a non-bijective epimorphism of solutions $f\colon(X,r)\to(Y,s)$ with $|Y|>1$. By Lemma 3.2, $\mathbb{D}_{f}$ is an ideal of $\mathcal{S}=\mathcal{S}(X,r)$ such that $\sim_{\mathbb{D}_{f}}$ is a refinement of $\sim_{f}$. Now, $\mathcal{D}_{f}$ is not the trivial ideal as $f$ is not bijective. Hence, by the assumptions, $\mathcal{D}\subseteq\mathcal{D}_{f}$ and thus $\mathbb{D}\subseteq\mathbb{D}_{f}$ and also $\mathbb{D}_{f}$ acts transitively on $X$. So, $x\sim_{\mathbb{D}_{f}}y$ for all $x,y\in X$. Therefore, $x\sim_{f}y$ for all $x,y\in X$, a contradiction. ∎
We now formulate the main result in what appears at first sight a more general context. However the following reasoning shows this actually is not the case. Let $B$ be a finite skew left brace and let $(B,r_{B})$ be its associated solution. Hence we have the associated skew left brace $\mathcal{S}(B)=\mathcal{S}(B,r_{B})$ that acts on the set $B$ (see Lemma 3.1) via the group $(B,+)\rtimes(B,\circ)$. So for any $x\in B$, $a,b\in B$ the element $(a,b)$ acts on $x$ as follows $\sigma_{a}^{-1}\lambda_{b}(x)$. If $X$ is a subset of $B$ and $I$ is an ideal of $B$ then we say that $I$ acts on $X$ if $X$ is invariant under the action of the subgroup $I\rtimes I$ of $(B,+)\rtimes(B,\circ)$.
Assume $X$ is a subset of $B$ that is an additive set of generators of $(B,+)$. Assume furthermore that $r_{B}$ restricts to a solution $r=r_{B}|_{X\times X}$ on $X$ (i.e., $X$ is invariant under the action of $\mathcal{S}(B,r_{B})$). Hence, we obtain natural morphisms of skew left braces
$G(X,r)\to G(B,r_{B})\to B\to\mathcal{G}(B,r_{B})\to\mathcal{G}(X,r).$
Since the composition of all these maps is the natural epimorphism $\operatorname{v}\colon G(X,r)\to\mathcal{G}(X,r)$, the composed map $B\to\mathcal{G}(B,r_{B})\to\mathcal{G}(X,r)$ is an epimorphism. If also $(X,r)$ is not a Lyubashenko solution then there exist distinct $x,y\in X$ with different images in $\mathcal{G}(X,r)$. Hence if the ideal $I$ generated by all $x-y$, with $x,y\in X\subseteq B$ is the smallest non-zero ideal of $B$ then it follows that $B\cong\mathcal{G}(B,r_{B})\cong\mathcal{G}(X,r)$.
So we obtain the following result.
Theorem 3.5.
Let $(X,r)$ be finite non-degenerate bijective solution of the Yang–Baxter equation that is not of Lyubashenko type. Then $(X,r)$ is simple if and only if $(X,r)$ is embedded into the associated solution $(B,r_{B})$ of a finite skew left brace $B$ such that $B$ is additively generated by $X$, the ideal $I$ generated by the set $\{x-y:x,y\in X\}$ is the smallest non-zero ideal of $B$, and the action of $I$ as an ideal of $B$ on $X$ is transitive.
Note that if $B$ is simple in the previous theorem then the conditions are equivalent with $B$ is additively generated by $X$ and the action of the group $\langle\sigma_{a}\lambda_{b}:a,b\in B\rangle$ on $X$ is transitive. It is worth mentioning that in [20, Remark 5.3] examples of finite skew simple left braces of abelian type $B$ are given (so the additive structure is commutative) so that for every orbit $X$ of any element under the action of the lambda map, $B\neq\langle X\rangle_{+}$.
Assume $(B,+,\circ)$ and $X$ are as in Theorem 3.5 and $I\neq B$. Then $B/I$ is a trivial skew left brace of cyclic type. In particular, $B^{2}=B*B\subseteq I$. As $B^{2}$ always is an ideal, and, by assumption $I$ is the unique minimal non-zero ideal of $B$, it follows that either $B^{2}=0$ (i.e., all $\lambda_{a}(b)-b=a*b=0$ and thus all $\lambda_{a}=\operatorname{id}$ and thus $(X,r)$ is a left derived solution) or $B^{2}=I$. In the latter case, $B/B^{2}$ is a trivial skew left brace of cyclic type and $B^{2}$ acts transitively on $X$.
If $B^{2}=0$, then $B$ is a trivial skew left brace and $(X,r)$ is a left derived solution. First note that $B$ cannot be of abelian type if $|X|>2$ (otherwise, the solution is the twist map and it is simple if and only if $|X|\leqslant 2$). Since by Theorem 3.4 $(X,r)$ is irretractable, it follows that $Z(B)$ is trivial (indeed, if $a\in Z(B)$ and $y\in X$, then $\sigma_{a}(y)\in X$ and $\sigma_{a}(y)=y$ hence $a=0$). Moreover, the unique minimal non-zero ideal $I$ needs to coincide with the derived subgroup $[B,B]$. This follows from the fact that $X$ is an orbit and so all generating elements in $I$ are commutators. Indeed, let $x-y$ be a generator of $I$, since $X$ is an orbit there exists $a\in B$ such that $\sigma_{a}(x)=y$, then $x-y=x+a-x-a=[x,a]$. Now, if $[B,B]\neq B$, since $X$ is a conjugacy class generating $B$ then $B/[B,B]$ needs to be cyclic. Note that this recovers the classification of simple quandle given by Joyce in [33] and which is stated in part two of the following corollary. We also mention that Andruskiewitsch and Graña, refined Joyce’s results in [1], giving a further characterization of groups with a unique minimal normal subgroup.
Hence we have shown the following reformulation.
Corollary 3.6.
A finite non-degenerate bijective solution $(X,r)$ of the Yang–Baxter equation is simple if and only if it satisfies one of the following types:
  1. (1)
    $(X,r)$ is a simple Lyubashenko solution,
  2. (2)
    $(X,r)$ is a left derived solution that is embedded into a conjugation quandle $(B,r_{B})$ defined by a finite group $B$ satisfying the following properties:
    • $X$ is a conjugacy class that generates $B$,
    • the derived subgroup $[B,B]$ is the smallest non-zero normal subgroup of $B$,
    • $B/[B,B]$ is a cyclic group,
    • $Z(B)$ is trivial.
  3. (3)
    $(X,r)$ is embedded into the associated solution $(B,r_{B})$ of a finite skew left brace $B$ with the following properties:
    • $B$ is additively generated by $X$,
    • the ideal $B^{2}$ is the smallest non-zero ideal of $B$,
    • $B/B^{2}$ is a trivial skew left brace of cyclic type,
    • the action of the ideal $B^{2}$ of $B$ on $X$ is transitive.
Remark 3.7.
Clearly, since there are finite simple non-degenerate involutive solutions $(X,r)$ of cardinality larger than $2$, the left derived solution $(X,r^{\prime})$ does not have to be simple if $(X,r)$ is simple. The converse however holds, as already noticed in [1].
Indeed, assume $f\colon(X,r)\to(Y,s)$ is an epimorphism of finite non-degenerate bijective solutions. Then one easily verifies that it induces an epimorphism $(X,r^{\prime})\to(Y,s^{\prime})$ between the left derived solution $(X,r^{\prime})$ of $(X,r)$ and the left derived solution $(Y,s^{\prime})$ of $(Y,s)$. If $(X,r^{\prime})$ is a simple solution, then we get that $|Y|=|X|$ (and thus $f$ is bijective) or $|Y|=1$. So a finite non-degenerate bijective solution $(X,r)$ is simple, if its derived solution $(X,r^{\prime})$ is simple.
4. Lyubashenko solutions
We describe the simple finite non-degenerate solutions that are of Lyubashenko type.
In [27] Etingof, Guralnick, and Soloviev have described the indecomposable non-degenerate bijective solutions of prime order. We now show that Lyubashenko solutions that are simple have to be of prime order and indecomposable.
Proposition 4.1.
Assume $(X,r)$ is a finite non-degenerate Lyubashenko solution with $|X|>1$, i.e., $r(x,y)=(\lambda(y),\rho(x))$ for some commuting permutations $\lambda,\rho\in\operatorname{Sym}(X)$. Then $(X,r)$ is simple if and only if $|X|=p$ for some prime $p$ and the group $H=\langle\lambda,\rho\rangle$ is cyclic of order $p$. In other words, identifying $X$ with $\mathbb{Z}_{p}$, such solutions are isomorphic with solutions of the form $\mathbb{Z}_{p}^{2}\to\mathbb{Z}_{p}^{2}\colon(x,y)\mapsto(x+a,y+b)$ with $a,b\in\mathbb{Z}_{p}$ and $(a,b)\neq(0,0)$. Actually we can take $a=1$ or $b=1$.
Proof.
To prove the necessity, assume $(X,r)$ is simple. Clearly, the result is obvious if $|X|=2$. So, we assume that $|X|>2$. Moreover, the derived solution is also of Lyubashenko type, i.e., $(x,y)\mapsto(y,\sigma(x))$, where $\sigma=\lambda\rho$. Hence $\mathcal{S}=\langle\sigma\rangle\rtimes\langle(\sigma^{-1},\lambda)\rangle$ and the action of $\mathcal{S}$ and $H$ on $X$ is the same.
Since $(X,r)$ is simple, it also is indecomposable. Hence, the action of the group $H$ and so also $\mathcal{S}$ is transitive on $X$.
Assume now that $\lambda$ or $\rho$ has a fixed point. By symmetry, we may assume that $\lambda(x)=x$ for some $x\in X$. Then, by transitivity, we get $X=Hx=\{\rho^{i}(x):i\geqslant 1\}$ and thus $\lambda(\rho^{i}(x))=\rho^{i}(\lambda(x))=\rho^{i}(x)$ implies $\lambda=\operatorname{id}$. Hence, $H=\langle\rho\rangle$ and if $\rho$ were not a full cycle on $X$ then again we would obtain a non-trivial surjective morphism $f\colon(X,r)\to(Y,s)$, constructed as in the previous paragraph. Hence $\rho$ must be a full cycle. Moreover, we claim that $|X|=p$ for some prime $p$. Indeed, otherwise $|X|=nm$ for some $n,m>1$ and then considering the set of orbits $Y=X/K$ with respect to the action of $K=\langle\rho^{n}\rangle\subseteq H$ on $X$ (note that $|Y|=n$ as $\rho^{n}$ is a product of $n$ disjoint cycles of length $m$ each) we would get again a surjective morphism $f\colon(X,r)\to(Y,s)$ with $1<|Y|<|X|$, a contradiction.
Finally, assume that both $\lambda$ and $\rho$ have no fixed points. Write $\lambda=\lambda_{1}\dotsm\lambda_{n}$, where the permutations $\lambda_{1},\dotsc,\lambda_{n}\in\operatorname{Sym}(X)$ are disjoint and such that each $\lambda_{i}=\gamma_{i1}\dotsm\gamma_{id_{i}}$ is a product of $d_{i}$ disjoint cycles $\gamma_{i1},\dotsc,\gamma_{id_{i}}$ of length $l_{i}$; we assume that $l_{1}<\dotsb<l_{n}$; as $\lambda$ has no fixed points, it follows that $l_{1}>1$. Since $\rho$ commutes with $\lambda$, it is well-known that
$\rho=\prod_{i=1}^{n}\gamma_{i1}^{k_{i1}}\dotsm\gamma_{id_{i}}^{k_{id_{i}}}\Theta_{i},$
where $1\leqslant k_{ij}\leqslant l_{i}$ for $1\leqslant i\leqslant n$ and $1\leqslant j\leqslant d_{i}$, and the permutation $\Theta_{i}\in\operatorname{Sym}(X)$ is determined by $\theta_{i}\in S_{d_{i}}$ in the following way: if $\gamma_{ij}=(x_{ij}^{1}\dotsm x_{ij}^{l_{i}})$ then
$\Theta_{i}(x)=\begin{cases}x_{i\theta_{i}(j)}^{k}&\text{if }x=x_{ij}^{k}\text{for some }1\leqslant j\leqslant d_{i}\text{ and }1\leqslant k\leqslant l_{i},\\ x&\text{otherwise}.\end{cases}$
(Note that $\Theta_{i}$ is responsible for permutation of “blocks”.) If $n>1$ then we again obtain a non-trivial surjective morphism $f\colon(X,r)\to(Y,s)$, where $Y=\{x_{11}^{1},\dotsc,x_{n1}^{1}\}$ and $f(x_{ij}^{k})=x_{i1}^{1}$ (that is, we may think of $(Y,s)$ and $f$ as obtained by identifying, for each $i$, elements in the support of the permutation $\lambda_{i}$), a contradiction with the simplicity of $(X,r)$. Thus $n=1$ and $\lambda=\lambda_{1}$ is a product of disjoint cycles of equal length. To simplify the notation we may write $\gamma_{j}=\gamma_{1j}$ and $x_{j}^{k}=x_{1j}^{k}$. Hence $\lambda=\gamma_{1}\dotsm\gamma_{d}$, where $d=d_{1}$ and $\gamma_{1},\dotsc,\gamma_{d}\in\operatorname{Sym}(X)$ are cycles of length $l=l_{1}$. Then, as we know,
$\rho=\gamma_{1}^{k_{1}}\dotsm\gamma_{d}^{k_{d}}\Theta,$
where $\Theta(x_{j}^{k})=x_{\theta(j)}^{k}$ for some $\theta\in S_{d}$ provided $\gamma_{j}=(x_{j}^{1},\dotsc,x_{j}^{l})$. Now if $Y=\{x_{1}^{1},\dotsc,x_{d}^{1}\}$ and $f\colon X\to Y$ is defined as $f(x_{j}^{k})=x_{j}^{1}$ (that is, we may view $(Y,s)$ and $f$ as obtained by identification of elements in each $\lambda$-orbit) then $f\colon(X,r)\to(Y,s)$ is a non-trivial surjective morphism of solutions as long as $d>1$. Hence $d=1$ and thus $\lambda=\gamma$ (where $\gamma=\gamma_{1}$) is a full cycle, whereas $\rho=\gamma^{k}$ (where $k=k_{1}$) is also a full cycle. Again, as in the third paragraph of the proof, one easily shows that that $|X|=p$ for some prime $p$. This finishes the proof of the necessity. The converse easily is verified. ∎
One easily can construct examples of simple finite non-degenerate bijective (derived) solutions of prime order that are not of Lyubashenko type.
Example 4.2.
Let $X=\{1,2,3\}$ and define $\rho_{1}=(23)$, $\rho_{2}=(13)$ and $\rho_{3}=(12)$ (this is the conjugation action in $S_{3}$ if one identifies $X$ with the set of elements of order $2$ in $S_{3}$). Then the solution $r(x,y)=(y,\rho_{y}(x))$ is simple by Lemma 5.6. Note that since all $\lambda_{x}=\operatorname{id}$ we have that $\sigma_{x}=\rho_{x}$ for each $x\in X$. Hence, for this example, $\mathcal{G}=\{(\rho_{a}^{-1},\operatorname{id}):a\in A\}\cong S_{3}$ and $\mathcal{C}=\{\rho_{a}:a\in A\}\cong S_{3}$. As this solution is not of Lyubashenko type, simple solutions that are not involutive are more complex than the involutive simple cases, as for the latter there is only one such solution of prime order (see [26]) and it is of Lyubashenko type, i.e., of the form $r(x,y)=(\lambda(y),\lambda^{-1}(x))$ with $\lambda$ a full cycle on $X$. Clearly this example can be generalized by making the size of $X$ equal to any prime $p>2$.
5. Examples
In this section we give some examples as applications of the results obtained.
Recall that any finite group $(G,\circ)$ can be considered as a skew left brace $(G,\circ,\circ)$, the trivial skew left brace on the group $G$. Note that in this skew left brace $\lambda_{a}=\operatorname{id}$ and $\rho_{a}=\sigma_{a}$ is conjugation by $a$. Hence, the associated solution is $r_{G}(a,b)=(b,b^{-1}\circ a\circ b)$ is a left derived solution, called the left conjugation solution. As mentioned earlier, in such a trivial skew left brace the ideals are precisely the normal subgroups of $(G,\circ)$. Of course in general one has several skew left brace structures on $G$. However, as a consequence on Hopf–Galois structures, Byott proves in [9] that if $(G,\circ)$ is a non-abelian simple group then there are only two possible skew left brace structures on $G$, namely the trivial skew left brace $(G,\circ,\circ)$ and the almost trivial skew left brace $(G,\circ^{\operatorname{op}},\circ)$, where $\circ^{\operatorname{op}}$ is the opposite operation of $\circ$. The latter corresponds with the bijective solutions $(X,r)$ with all $\rho_{x}=\operatorname{id}$ and thus $\sigma_{x}=\lambda^{-1}_{x}$ for all $x\in X$. In this case the associated solution is $r_{G}(a,b)=(a\circ b\circ a^{-1},a)$, called the right conjugation solution.
Note that a finite non-abelian simple group $G$ is generated by any non-trivial conjugacy class.
Example 5.1.
Assume that $(X,r)$ is a finite non-degenerate bijective left derived solution that is not of Lyubashenko type. If the group $\mathcal{G}(X,r)$ is simple non-abelian, then $\mathcal{G}(X,r)$ has only two possible skew left brace structures, the trivial and the almost trivial. Only the former yields left derived solutions. Hence, from Corollary 3.6 and the discussion in Section 3, we obtain: $(X,r)$ is a simple solution if and only if $\{\sigma_{x}:x\in X\}$ is a conjugacy class in $\mathcal{G}(X,r)$, the map $v\colon X\to\mathcal{G}(X,r)\colon x\mapsto\sigma_{x}$ is injective and $r(x,y)=(v^{-1}(\sigma_{x}^{-1}\sigma_{y}\sigma_{x}^{-1}),x)$. In other words, the associated solution of a finite non-abelian simple group $G$ induces a simple solution on a subset $X$ if and only if $X$ is a conjugacy class of $G$.
Since conjugacy classes can have different cardinality we have constructed within a simple non-abelian group simple solutions of different cardinality.
A concrete example is the alternating group $A_{n}$ of degree $n\geqslant 5$ and $X$ the set of $3$-cycles. It is of cardinality $2\binom{n}{3}$. If $n=7$ we get $|X|=70=2\cdot 5\cdot 7$, an even square-free number ($n=5$ and $n=6$ give non-square-free numbers $20=2^{2}\cdot 5$ and $40=2^{3}\cdot 5$, respectively). This is in contrast with the involutive case (see [21]) where it is shown that in this case all simple involutive solutions have cardinality, which is not a square free number. This follows from a much more general result: a finite indecomposable non-degenerate involutive set-theoretic solution of the Yang–Baxter equation on a set $X$ of square-free cardinality is a multipermutation solution.
Remark 5.2.
[…]
Byott mentions that there are only three types of groups of order $n=p^{p}q$. The Sylow $p$-subgroup $P$ or the Sylow $q$-subgroup $Q$ has to be normal (and thus invariant) and the three groups are respectively the direct product $P\times Q$ and the semi-direct products $P\rtimes Q$ and $Q\rtimes P$, of which he utilizes the latter two. A crucial property of those last two groups is that neither contains elements of order $pq$ and the latter does not contain a normal subgroup of order $p^{i}$ for some $i$. This is well-known and follows from the assumptions on $p$ and $q$ as $\operatorname{GL}_{p}(\mathbb{F}_{p})$ will be the linear group of the smallest dimension containing an element of order $q$. To yield a skew left brace structure $B$ one needs that $P$ is normal and elementary abelian (hence one may identify $P$ with the additive group of the $\mathbb{F}_{p}$-vector space $V=\mathbb{F}_{p}^{p}$), and the group has normal subgroups of order $1$, $p^{p}$ and $n$ and no subgroups of order $p^{i}q$ for $1\leqslant i<p$. It is then shown that $(B,+)\cong P\rtimes Q$ and $(B,\circ)\cong Q\rtimes P$.
As an application of Theorem 3.4, we show that these simple skew left braces yield simple non-degenerate bijective solutions of size $p^{p}(q-1)$. To do so we recall from [10] an explicit construction. Let $J\in\operatorname{GL}_{p}(\mathbb{F}_{p})$ be a Jordan matrix with eigenvalue $1$. Then, there exists a matrix $M\in\operatorname{GL}_{p}(\mathbb{F}_{p})$ of order $q$ with $JMJ^{-1}=M^{p}$. Put $V=\mathbb{F}_{p}^{p}$ and consider the natural left action of the group $\langle M\rangle$ on $V$ and the associated semi-direct product group
$B=V\rtimes\langle M\rangle=\{(v,M^{i}):v\in V\text{ and }0\leqslant i<q\},$
written additively. Clearly $(B,+)\cong C_{p}^{p}\rtimes C_{q}$. Then, according to [10] one can construct a regular subgroup $(B,\circ)$ of the holomorph $\operatorname{Hol}(B,+)=(B,+)\rtimes\operatorname{Aut}(B,+)$ with generating elements:
$x$ $=((0,M),\operatorname{id}),$
[…]
$z$ $=((e_{p},I),\operatorname{conj}(-e_{p},J)),$
where $V_{0}=\mathbb{F}_{p}e_{1}\oplus\dotsb\oplus\mathbb{F}_{p}e_{p-1}$, $\{e_{1},\dotsc,e_{p}\}$ is the standard basis of $V$, and $\operatorname{conj}(b)\in\operatorname{Aut}(B,+)$ denotes the conjugation by $b\in B$. Then one can prove that the skew left brace $(B,+,\circ)$ is simple. Denote by $Q$ the cyclic group of order $q$ generated by $x$. We claim that the set
$X=\bigcup_{b\in B}(b+Q-b)\setminus\{1\}$
generate $(B,+)$ and $B$ acts transitively on $X$. We first prove the latter. Note that $X$ coincides with the set of elements of order $q$ in $(B,+)$. Hence, $X$ is closed under the action of $B$, as all $\lambda_{b}$ and $\sigma_{b}$ maps are automorphisms of $(B,+)$. To show that the action is transitive, it is sufficient to prove that we can reach, under the action, any element of $Q$ starting from $x\in Q$, as $\sigma$ coincides with conjugation. For this, one first notices that
$z+\lambda_{z}(x)-z$ $=z+(-e_{p}+JMJ^{-1}e_{p},JMJ^{-1})-z$
$=(e_{p}-e_{p}+JMJ^{-1}e_{p}-JMJ^{-1}e_{p},M^{p})=(0,M^{p})=px.$
As $p$ is invertible modulo $q$, one then can inductively reach all elements in $Q\setminus\{1\}$. Rests to show that the set $X$ additively generates $B$. Hence, this shows the transitivity on $X$. Hence, $X$ coincides with all elements of $B$ that are not of $p$ power order, i.e., the complement of $V\rtimes\{I\}$. Thus,
$X=\{((v,M^{i}),\lambda_{(v,M^{i})}):v\in V\text{ and }0<i<q\}.$
As $(v,I)=(v,M)-(0,M)$, the claim follows and $X$ indeed generates $(B,+)$. Hence, by Corollary 3.6, the solution restricted to $X$ is a simple solution of size $p^{p}(q-1)$.
Also within some non-simple groups we easily can construct simple solutions.
The following example appears in [1] as an example of a simple quandle.
Example 5.3.
Let $S_{n}$ be the symmetric group of degree $n$ and consider it as an trivial skew left brace $(S_{n},\circ,\circ)$ with its associated left conjugation solution. So each $\rho_{g}$ is conjugation, i.e., $\rho_{g}(h)=g^{-1}\circ h\circ g$ and each $\lambda_{g}=\operatorname{id}$. Let $X$ be the conjugacy class consisting of all transpositions in $S_{n}$. If $n\geqslant 5$, the only ideal of this skew left brace is $A_{n}$. Hence it is the smallest non-zero ideal of $S_{n}$. Since all elements of the form $g\circ h^{-1}$ lie in $A_{n}$, with $g,h\in X$, and because $A_{n}$ is generated as a group by $X$, it follows from Theorem 3.5 that the solution $(X,r)$ is a simple with $r(g,h)=(h,h^{-1}\circ g\circ h)$. So, we have shown the existence of simple solutions of cardinality $|X|=\binom{n}{2}$ for any integer $n\geqslant 5$. If for example $n=6$ then $|X|=15=3\cdot 5$ and again we have shown the existence of simple solutions of square-free cardinality that are not prime numbers.
Another example is the following.
Example 5.4.
Let $(D_{2p},\circ)$ be the dihedral group of order $2p$ and again equip it with the almost trivial skew left brace structure $(D_{2p},\circ^{\operatorname{op}},\circ)$. The set $X$ of elements of order $2$ is a conjugacy class and the only normal subgroup of $D_{2p}$ is the cyclic subgroup of index $2$. Hence we obtain a simple solution $(X,r)$ of cardinality $p$.
Example 5.5.
All indecomposable non-degenerate bijective solutions of prime order are simple solutions.
This can easily be proven. First, as in [12] for involutive solutions, we say that a morphism $f\colon(X,r)\to(Y,s)$ of solutions is a covering map if it is surjective and all the fibres $f^{-1}(y)$ for $y\in Y$ have the same cardinality. A non-degenerate bijective solution $(X,r)$ is said to be transitive if $\mathcal{S}(X,r)$ acts transitively on $X$. The following lemma was proven for non-degenerate involutive solutions in [12, Lemma 1]. The proof remains the same. For completeness’ sake we include a proof.
Lemma 5.6.
Assume $(X,r)$ is an indecomposable finite non-degenerate bijective solution of the Yang–Baxter equation. If $f\colon(X,r)\to(Y,s)$ is an epimorphism of solutions then $f$ is a covering map. In particular, if $|X|$ is prime then $(X,r)$ is a simple solution.
Proof.
By assumption $(X,r)$ is indecomposable and so $\mathcal{S}=\mathcal{S}(X,r)$ acts transitively on $X$. Now, if $x\in f^{-1}(y)$ with $y\in Y$ then, for any $\alpha=(\sigma_{a},\operatorname{v}_{b})\in\mathcal{S}$, we have
$f(\alpha\cdot x)=f(\sigma_{a}\lambda_{b}(x))=\sigma_{f(a)}\lambda_{f(b)}(f(x))=\sigma_{f(a)}\lambda_{f(b)}(y).$
So, $\alpha\cdot x\in f^{-1}(\sigma_{f(a)}\lambda_{f(b)}(y))$. Hence, if $x,x^{\prime}\in f^{-1}(y)$ are distinct then $\alpha\cdot x,\alpha\cdot x^{\prime}\in f^{-1}(\sigma_{f(a)}\lambda_{f(b)}(y))$ and $\alpha\cdot x\neq\alpha\cdot x^{\prime}$. Therefore, it follows that $|f^{-1}(y)|\leqslant|f^{-1}(\sigma_{f(a)}\lambda_{f(b)}(y))|$. Clearly, because $f$ is an epimorphism, the group $\mathcal{S}(Y,s)$ acts transitively on $Y$. Thus all fibres $f^{-1}(y)$ have the same cardinality, and indeed $f$ is a covering map.
To prove the second statement, assume $|X|$ is a prime number. From the first part it follows $|Y|$ is a divisor of $|X|$. Hence, $|Y|=|X|$ or $|Y|=1$ and thus $(X,r)$ is a simple solution. ∎
In [27] Etingof, Guralnick and Soloviev have described finite indecomposable non-degenerate bijective solutions $(X,r)$ of the Yang–Baxter equation that are of prime order. It turns that then $X$ can be equipped with an abelian group structure $(X,+)$ (and thus we may assume that $X=\mathbb{Z}_{p}$) such that the solution $(X,r)$ is affine, i.e., it is of the form
$r(x,y)=(\alpha(x)+\beta(y)+a,\gamma(x)+\delta(y)+b)$
for some $\alpha,\beta,\gamma,\delta\in\operatorname{End}(X,+)$ and $a,b\in X$. They are precisely the solutions of the following type built on triples $a,b,c\in X$ with $a,b\neq 0$:
  1. (1)
    $(x,y)\mapsto(a^{-1}y+(1-a^{-1}b)x+h,bx-ac)$,
  2. (2)
    $(x,y)\mapsto(a^{-1}y+c,bx+(1-a^{-1}b)y-bc)$,
  3. (3)
    $(x,y)\mapsto(y+p,x+q)$, with $(0,0)\neq(p,q)\in X\times X$.
The solutions of the third type are precisely the Lyubashenko examples obtained in Proposition 4.1.
It is interesting to note that all these solutions satisfy $\lambda_{x}=\lambda_{y}$ or $\rho_{x}=\rho_{y}$ for all $x,y\in X$, and thus are of the type investigated in [22, Section 8], where it is shown that the Yang–Baxter algebras corresponding to such solutions always are left, respectively right, Noetherian.
Acknowledgement
The first named author is supported by the Engineering and Physical Sciences Research Council [grant number EP/V005995/1] and the Fonds voor Wetenschappelijk Onderzoek (Flanders) – Krediet voor wetenschappelijk verblijf in Vlaanderen grant V512223N. The third author is supported by National Science Centre grant 2020/39/D/ST1/01852 (Poland). The fourth author is supported by Fonds Wetenschappelijk Onderzoek Vlaanderen (grant 1229724N).